OpenStudy (anonymous):

Find the average rate of change for the given function from x = 1 to x = 2.

OpenStudy (anonymous):

-2 \[-\frac{ 1 }{ 2 }\] \[\frac{ 1 }{ 2 }\] 2

OpenStudy (anonymous):

I don't get how to do this @ganeshie8 @satellite73

OpenStudy (anonymous):

they are asking for the slope

OpenStudy (anonymous):

the "average rate of change" of a line is a constant, it is the slope it has nothing to do with 1 or 2, it is the same everywhere

OpenStudy (anonymous):

so would it be 2?

OpenStudy (anonymous):

6/3 = 2/1 = 2, 4/2 = 2/1 = 2 and so on

OpenStudy (anonymous):

Yep 2 was right, thanks for the help